# Is Sora a World Simulator? A Comprehensive Survey on General World Models and Beyond

### v1

@@ 2.3.3 Towards World Models
Dataset Year Text Domain ##\##Video Avg Video len Avg text len Resolution
@@ 3.2 Neural Driving Simulator
2D GAIA-1 [91] Wayve [91] GPT VQ-VAE Video Diffusion Decoder Img, Text, Act Img
@@
3D UniWorld [151] nuScenes [25] Transformer BEV Encoder Task Specific Decoder Img Img, Occ
@@ 4.3 Commonly Used Benchmarks
TABLE V: Game scores and human-normalized scores (HNS in %) of world-model-based game agents on the 26 games in the the Atari[14] 100k benchmark[105]. The last two columns are the scores achieved by DeepMind human gamers and the scores achieved by a random agent. The highest score of each row is bolded.
SimPLe [105] DreamerV3 [74] IRIS [146] TWM [175] STORM [247] HarmonyDreamer [141] Human Random
Alien 616.9 (5.6%) 959.0 (10.6%) 420.0 (2.8%) 674.6 (6.5%) 984.0 (11.0%) 890.0 (9.6%) 7127.7 (100.0%) 227.8 (0.0%)
Amidar 74.3 (4.0%) 139.0 (7.8%) 143.0 (8.0%) 121.8 (6.8%) 205.0 (11.6%) 141.0 (7.9%) 1719.5 (100.0%) 5.8 (0.0%)
Assault 527.2 (58.7%) 706.0 (93.1%) 1524.4 (250.6%) 682.6 (88.6%) 801.0 (111.4%) 1003.0 (150.2%) 742.0 (100.0%) 222.4 (0.0%)
Asterix 1128.3 (11.1%) 932.0 (8.7%) 853.6 (7.8%) 1116.6 (10.9%) 1028.0 (9.9%) 1140.0 (11.2%) 8503.3 (100.0%) 210.0 (0.0%)
BankHeist 34.2 (2.7%) 649.0 (85.9%) 53.1 (5.3%) 466.7 (61.2%) 641.0 (84.8%) 1069.0 (142.8%) 753.1 (100.0%) 14.2 (0.0%)
BattleZone 4031.2 (4.8%) 12250.0 (28.4%) 13074.0 (30.8%) 5068.0 (7.8%) 13540.0 (32.1%) 16456.0 (40.5%) 37187.5 (100.0%) 2360.0 (0.0%)
Boxing 7.8 (64.2%) 78.0 (649.2%) 70.1 (583.3%) 77.5 (645.0%) 80.0 (665.8%) 80.0 (665.8%) 12.1 (100.0%) 0.1 (0.0%)
Breakout 16.4 (51.0%) 31.0 (101.7%) 83.7 (284.7%) 20.0 (63.5%) 16.0 (49.7%) 53.0 (178.1%) 30.5 (100.0%) 1.7 (0.0%)
ChopperCommand 979.4 (2.6%) 420.0 (-5.9%) 1565.0 (11.5%) 1697.4 (13.5%) 1888.0 (16.4%) 1510.0 (10.6%) 7387.8 (100.0%) 811.0 (0.0%)
CrazyClimber 62583.6 (206.8%) 97190.0 (345.0%) 59324.2 (193.8%) 71820.4 (243.7%) 66776.0 (223.5%) 82739.0 (287.3%) 35829.4 (100.0%) 10780.5 (0.0%)
DemonAttack 208.1 (3.1%) 303.0 (8.3%) 2034.4 (103.5%) 350.2 (10.9%) 165.0 (0.7%) 203.0 (2.8%) 1971.0 (100.0%) 152.1 (0.0%)
Freeway 16.7 (56.4%) 0.0 (0.0%) 31.1 (105.1%) 24.3 (82.1%) 34.0 (114.9%) 0.0 (0.0%) 29.6 (100.0%) 0.0 (0.0%)
Frostbite 236.9 (4.0%) 909.0 (19.8%) 259.1 (4.5%) 1475.6 (33.0%) 1316.0 (29.3%) 679.0 (14.4%) 4334.7 (100.0%) 65.2 (0.0%)
Gopher 596.8 (15.7%) 3730.0 (161.1%) 2236.1 (91.8%) 1674.8 (65.8%) 8240.0 (370.4%) 13043.0 (593.3%) 2412.5 (100.0%) 257.6 (0.0%)
Hero 2656.6 (5.5%) 11161.0 (34.0%) 7037.4 (20.2%) 7254.0 (20.9%) 11044.0 (33.6%) 13378.0 (41.4%) 30826.4 (100.0%) 1027.0 (0.0%)
Jamesbond 100.5 (26.1%) 445.0 (151.9%) 462.7 (158.4%) 362.4 (121.8%) 509.0 (175.3%) 317.0 (105.2%) 302.8 (100.0%) 29.0 (0.0%)
Kangaroo 51.2 (0.0%) 4098.0 (135.6%) 838.2 (26.4%) 1240.0 (39.8%) 4208.0 (139.3%) 5118.0 (169.8%) 3035.0 (100.0%) 52.0 (0.0%)
Krull 2204.8 (56.8%) 7782.0 (579.3%) 6616.4 (470.1%) 6349.2 (445.1%) 8413.0 (638.4%) 7754.0 (576.7%) 2665.5 (100.0%) 1598.0 (0.0%)
KungFuMaster 14862.5 (65.0%) 21420.0 (94.1%) 21759.8 (95.7%) 24554.6 (108.1%) 26182.0 (115.3%) 22274.0 (97.9%) 22736.3 (100.0%) 258.5 (0.0%)
MsPacman 1480.0 (17.6%) 1327.0 (15.3%) 999.1 (10.4%) 1588.4 (19.3%) 2673.0 (35.6%) 1681.0 (20.7%) 6951.6 (100.0%) 307.3 (0.0%)
Pong 12.8 (94.9%) 18.0 (109.6%) 14.6 (100.0%) 18.8 (111.9%) 11.0 (89.8%) 19.0 (112.5%) 14.6 (100.0%) -20.7 (0.0%)
PrivateEye 35.0 (0.0%) 882.0 (1.2%) 100.0 (0.1%) 86.6 (0.1%) 7781.0 (11.2%) 2932.0 (4.2%) 69571.3 (100.0%) 24.9 (0.0%)
Qbert 1288.8 (8.5%) 3405.0 (24.4%) 745.7 (4.4%) 3330.8 (23.8%) 4522.0 (32.8%) 3933.0 (28.4%) 13455.0 (100.0%) 163.9 (0.0%)
RoadRunner 5640.6 (71.9%) 15565.0 (198.6%) 9614.6 (122.6%) 9109.0 (116.1%) 17564.0 (224.1%) 14646.0 (186.8%) 7845.0 (100.0%) 11.5 (0.0%)
Seaquest 683.3 (1.5%) 618.0 (1.3%) 661.3 (1.4%) 774.4 (1.7%) 525.0 (1.1%) 665.0 (1.4%) 42054.7 (100.0%) 68.4 (0.0%)
UpNDown 3350.3 (25.2%) 7667.0 (63.9%) 3546.2 (27.0%) 15981.7 (138.4%) 7985.0 (66.8%) 10874.0 (92.7%) 11693.2 (100.0%) 533.4 (0.0%)
Avg. HNS 33.2% 112.4% 104.6% 95.6% 126.7% 136.6% 100.0% 0.0%
@@
Dreamer [72] DreamerV2 [73] Dreaming [157] DreamingV2 [158] DreamerPro [41]

### v2

@@ 2.3.3 Towards World Models
Dataset Year Text Domain #\#Video Avg Video len Avg text len Resolution
@@ 3.2 Neural Driving Simulator
2D
GAIA-1 [28] Wayve [28] GPT VQ-VAE Video Diffusion Decoder Img, Text, Act Img
@@
3D
UniWorld [189] nuScenes [185] Transformer BEV Encoder Task Specific Decoder Img Img, Occ
@@ 4.3 Commonly Used Benchmarks
TABLE V: Human-normalized scores (HNS in %) of world-model-based agents on 26 games in the Atari100k benchmark[223]. The highest score of each row is bolded.
SimPLe
[223]
DreamerV3
[35]
IRIS
[215]
TWM
[212]
STORM
[211]
Harmony
Dreamer
[239]
DIAMOND
[globerson2024diamond]
Alien 5.6 10.6 2.8 6.5 11.0 9.6 7.5
Amidar 4.0 7.8 8.0 6.8 11.6 7.9 12.8
Assault 58.7 93.1 250.6 88.6 111.4 150.2 251.0
Asterix 11.1 8.7 7.8 10.9 9.9 11.2 42.1
BankHeist 2.7 85.9 5.3 61.2 84.8 142.8 0.7
BattleZone 4.8 28.4 30.8 7.8 32.1 40.5 6.7
Boxing 64.2 649.2 583.3 645.0 665.8 665.8 723.3
Breakout 51.0 101.7 284.7 63.5 49.7 178.1 454.2
Chopper. 2.6 -5.9 11.5 13.5 16.4 10.6 8.5
CrazyClimber 206.8 345.0 193.8 243.7 223.5 287.3 352.9
DemonAttack 3.1 8.3 103.5 10.9 0.7 2.8 7.5
Freeway 56.4 0.0 105.1 82.1 114.9 0.0 112.5
Frostbite 4.0 19.8 4.5 33.0 29.3 14.4 4.9
Gopher 15.7 161.1 91.8 65.8 370.4 593.3 261.7
Hero 5.5 34.0 20.2 20.9 33.6 41.4 15.4
Jamesbond 26.1 151.9 158.4 121.8 175.3 105.2 145.5
Kangaroo 0.0 135.6 26.4 39.8 139.3 169.8 178.7
Krull 56.8 579.3 470.1 445.1 638.4 576.7 656.9
KungFuMaster 65.0 94.1 95.7 108.1 115.3 97.9 82.1
MsPacman 17.6 15.3 10.4 19.3 35.6 20.7 24.8
Pong 94.9 109.6 100.0 111.9 89.8 112.5 116.4
PrivateEye 0.0 1.2 0.1 0.1 11.2 4.2 0.1
Qbert 8.5 24.4 4.4 23.8 32.8 28.4 32.6
RoadRunner 71.9 198.6 122.6 116.1 224.1 186.8 263.8
Seaquest 1.5 1.3 1.4 1.7 1.1 1.4 1.1
UpNDown 25.2 63.9 27.0 138.4 66.8 92.7 29.8
Avg. HNS 33.2 112.4 104.6 95.6 126.7 136.6 145.9
@@
Dreamer
[33]
DreamerV2
[34]
Dreaming
[232]
DreamingV2
[205]
DreamerPro
[204]